# CCFExp: Facial Image Synthesis with Cycle Cross-Fusion Diffusion Model for Facial Paralysis Individuals

### v1

@@ Abstract
Facial paralysis is a debilitating condition that affects the movement of facial muscles, leading to a significant loss of facial expressions. Currently, the diagnosis of facial paralysis remains a challenging task, often relying heavily on the subjective judgment and experience of clinicians, which can introduce variability and uncertainty in the assessment process. One promising application in real-life situations is the automatic estimation of facial paralysis. However, the scarcity of facial paralysis datasets limits the development of robust machine learning models for automated diagnosis and therapeutic interventions. To this end, this study aims to synthesize a high-quality facial paralysis dataset to address this gap, enabling more accurate and efficient algorithm training. Specifically, a novel Cycle Cross-Fusion Expression Generative Model (CCFExp) based on the diffusion model is proposed to combine different features of facial information and enhance the visual details of facial appearance and texture in facial regions, thus creating synthetic facial images that accurately represent various degrees and types of facial paralysis. We have qualitatively and quantitatively evaluated the proposed method on the commonly used public clinical datasets of facial paralysis to demonstrate its effectiveness. Experimental results indicate that the proposed method surpasses state-of-the-art methods, generating more realistic facial images and maintaining identity consistency. Code Link
@@ 1 Introduction
Facial palsy, which can result in temporary or permanent weakness or paralysis of one side of the face, typically presents as an acute, unilateral facial nerve weakness or paralysis that develops rapidly, often within 72 hours, with an unknown cause, affecting approximately 23 out of every 100,000 people annually [15]. Beyond the physical discomfort, patients with facial palsy often experience significant psychological distress [21]. Accurate and timely diagnosis is crucial for effective treatment; however, the variability in clinical presentation poses significant challenges. The severity estimation of facial palsy remains a largely subjective process [15].
@@
In recent years, machine learning (ML) has shown promise in aiding the diagnosis of facial paralysis by analyzing facial features and expressions [62, 1, 14, 15]. However, the development of reliable ML models is hindered by the scarcity of comprehensive, annotated datasets that accurately represent the diversity and severity of facial paralysis. Obtaining high-quality facial images of facial palsy patients is a significant challenge due to several factors, including the need to protect patient privacy and the inherent sensitivity of medical data. Additionally, the rarity of the condition and the variability in its presentation make it difficult to gather a diverse and representative dataset. Existing datasets [64, 17, 24] are often limited in size, scope, and variability, restricting the generalization ability and performance of ML algorithms.
@@
Several studies have explored the use of generative adversarial networks (GANs) [16] to generate facial paralysis face images [66, 50, 49]. Although [66] successfully utilizes GANs to simulate facial expressions by reconstructing facial images from extracted boundary information, the generated images often lack the fine details necessary for accurate medical diagnosis, especially when there is a significant disparity between the facial features of the patient and the face to be combined with the palsy features. Building on this, another study [50] introduces facial shape normalization to address the unnatural appearance of images resulting from shape differences between the patient and the model. While this method improves image realism, it still struggles with capturing complex facial expressions and fine details, limiting its effectiveness in more nuanced diagnostic scenarios. In addition, to leverages data augmentation, GANs are also used to improve the performance of convolutional neural networks (CNNs) in grading facial palsy [49]. This approach synthesizes additional face images with varying degrees of facial palsy to combat overfitting and enhance the model’s generalization capability. However, the synthesized images in this study primarily focus on only these five categories without adequately capturing the full range of facial expressions and subtle asymmetries inherent in real-world cases.
@@
In this paper, we aim to address the critical need for a comprehensive facial paralysis dataset by synthesizing a large-scale, high-quality dataset. The generator is designed to capture a wide range of facial paralysis conditions and able to generate face images with varying degrees of facial paralysis based on a given facial identity image, providing a valuable resource for training ML models with the potential to improve diagnostic accuracy and clinical outcomes.
@@
To this end, we propose a Cycle Cross-Fusion Expression Generative Model (CCFExp). CCFExp is a unique triple condition image generation model, conditioned on identity information of a given ID image, expression information and landmark information of a given style image. It could generate a realistic image of the face in the ID image exhibiting the facial palsy characteristics of the face in the style image. CCFExp employs landmark features and the corresponding loss to ensure that the model could fully capture the nuanced facial deformations, particularly the detailed changes in facial features such as eyes, mouth, and overall asymmetry. When the model is fed multiple features simultaneously, it often encounters challenges such as feature redundancy and conflicting information, which can lead to degraded performance [35]. The redundancy among features can cause the model to overweight certain aspects, while conflicting signals may lead to inconsistent learning and prediction. To address these issues, CCFExp utilize a novel cross-fusion feature integration method. This approach effectively harmonizes the information from multiple features, enabling the model to better leverage the complementary aspects of each feature, thereby significantly improving the overall performance and robustness of the model. Training diffusion models [19, 53] on limited datasets often leads to inefficiencies in learning and suboptimal model performance due to the scarcity of data. To address these challenges, we focused on improving the training efficiency and overall performance of diffusion models under constrained data conditions by modifying the training strategy. Our approach optimizes the learning process and proposes the cycle diffusion framework enabling the model to achieve superior results with a limited dataset.
Our approach extends beyond the conventional five severity levels, incorporating a wider range of facial expressions and detailed deformations of facial features, ensuring that the subtle yet critical nuances of facial movement and deformation are accurately represented. This results in a more robust model capable of generalizing to various real-world cases, thereby enhancing its clinical applicability. Moreover, this work lays the foundation for future research in automated facial analysis and the synthesis of medical datasets.
@@
  • To address the shortage of medical datasets, our study pioneers the application of generative models in facial paralysis face synthesis while also proposing a new evaluation metric for conditionally generated objects, achieving novel SOTA results.
Figure 2: An overview of the forward process of CCFExp. a) The architecture of our feature extractors. We train the model using different images of the same patient, with one serving as the identity image $X_{ID}$ and the other as the facial palsy expression image $X_{0}$. The identity model extracts identity features from the identity image. The expression module (including $F_{Face}$ and $F_{ID}$) and $F_{LM}$ extract expression and landmark features from the facial palsy expression image, respectively. b) An illustration of the principle of the cross feature fusion strategy. We employ a method similar to cross-attention to facilitate information exchange among the three conditional features. After this interaction, the features are concatenated for further processing. c) A diagram of the noise predicting network. CCFExp utilizes a classic U-Net architecture combined with residual connections and cross-attention mechanisms to accurately predict the noise.
@@ 3 Method
In this section, we present our Cycle Cross-Fusion Expression Model (CCFExp) designed to synthesize high-quality facial paralysis images (see Fig. 2). First, we mention the necessary preliminary concepts about the diffusion models (sec 3.1). Next, we introduce the various feature extractors used in CCFExp which enhance the diffusion model’s ability to capture more facial details by accurately extracting multiple features (sec 3.2). In sec 3.3, we propose the cross feature fusion, which mitigates the model’s over-reliance on individual features and enhances its overall learning capability across all features by exchanging query vectors formed from different features. Finally, we devise the cycle training strategy that improves the training efficiency and performance of the diffusion model on limited datasets by modifying the training process and introducing a secondary loss function (sec 3.4).
@@ 3.1 Preliminary
$L=\mathbb{E}_{t,X_{t},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(X_{t},t)\|^{2}\right],$ (2)
@@ 3.2 Feature Extraction Module
ID Extractor
To accurately capture identity information from images, we use the IR-50 [9], a typical CNN, as the identity feature extractor. We utilize both the output features $f_{ID}\in\mathbb{R}^{512}$ from the final layer and the intermediate layer features $I_{ID}\in\mathbb{R}^{512\times 7\times 7}$of the model to enhance the quality of image generation. For $I_{ID}$, we first flatten and rearrange their dimensions to obtain $F_{ID}\in\mathbb{R}^{49\times 512}$ that facilitates subsequent processing. Then, we concatenate the two sets of features and add positional embeddings. In summary, the output of the identity feature extraction module $E_{ID}$ is formulated as follows:
$E_{ID}(X_{ID})=Concat[f_{ID},Flatten(I_{ID})]+P_{emb}.$ (3)
@@
Expression Extractor
CCFExp employs the deviation framework of DLN [70] to accurately capture the expression features. This module consists of two models: a facial feature extraction model $F_{Face}$ and an identity feature extraction model $F_{ID}$, sharing the same network structure. Based on the assumption that facial features can be divided into identity features and expression features, we use two separate models to extract expression and identity features from the same facial image. By subtracting the identity features from the facial features, we obtain expression features that are invariant to identity. The subtracted features are then passed through a convolution operation for channel alignment. Specifically,
$E_{Exp}(X_{0})=Conv[F_{Face}(X_{0})-F_{ID}(X_{0})].$ (4)
@@
Facial landmarks refer to key points on a face such as the corners of the eyes, the tip of the nose, and the edges of the lips that are used to represent and understand facial geometry, which could enhance facial generation by imposing constraints on critical facial regions. For the first time, we integrate facial landmark features with diffusion models, coupled with a specifically designed loss function, to generate realistic and detailed facial images of patients with facial palsy. In our CCFExp, we employ the MobileFaceNet [6] as the landmark detector $D_{LM}$. We remove its output layer to use it as a landmark feature extractor $E_{LM}$, which outputs a feature $F_{LM}\in\mathbb{R}^{49\times 512}$.
@@ 3.3 Cross-Fusion Module
The attention mechanism [61] has become a fundamental component in many state-of-the-art models, particularly for tasks requiring the integration of information from different sources or across different parts of a data sequence. At its core, the attention mechanism allows a model to weigh the importance of different elements within a sequence, dynamically focusing on the most relevant parts to improve decision-making or feature extraction. The cross-attention mechanism extends this concept to operate between two distinct sequences. This facilitates the fusion of information from different sources, making it particularly useful for tasks that require integrating features from multiple modalities or input streams:
[…]
where $Q_{1}$ is from the first sequence and $K_{2}$, $V_{2}$ are from the second sequence.
We aim for the model to fully utilize the information from identity, expression, and landmarks when predicting noise. Although directly concatenating the features from these three sources and injecting them into the model is a feasible approach, it does not yield satisfactory results. The generated images either lose the identity information from $X_{ID}$ or fail to retain the facial expression details from $X_{0}$. Building on the concept of the cross-attention mechanism, we perform mutual interaction among multiple features, where the query sequence from one feature guides the extraction of information from the key-value sequence of another feature.
The features from the extractors are first mapped into the query, key, and value vector spaces through a QKV encoder. Specifically,
@@
where $F$ indicates features, $W^{Q},W^{K},W^{V}$ represent the model’s learnable parameter matrixes. Similar to the multi-head attention mechanism [61], we utilize multiple sets of transformation matrices to obtain multiple vector spaces, aiming to capture different interaction patterns among features.
Subsequently, the identity features, expression features, and landmark features each have their own corresponding QKV vector spaces. Next, we perform vector interactions. Specifically, the query vectors from the identity features $Q_{ID}$ guide the extraction of information from the expression features, the query vectors from the expression features $Q_{E}$ guide the extraction of information from the landmark features, and the query vectors from the landmark features $Q_{LM}$ guide the extraction of information from the identity features. Through Eq. 5, we obtain the three interacted features, which are then concatenated together and subsequently fed into the noise prediction network.
Figure 3: An illustration of the cycle training strategy of CCFExp. The black unidirectional arrows represent the forward diffusion process, the blue and green double-arrow dashed lines indicate the calculation processes of the first and second losses, respectively. The red dotted points represent the landmark keypoints and the brown rectangular blocks represent the extracted identity features.
@@ 3.4 Cycle Diffusion Strategy
Building upon the training framework of the original diffusion model, we incorporated an additional step of secondary diffusion. Algorithm 1 shows the training procedure of CCFExp. After the original diffusion process, where the input image $X_{0}$ is noised to generate a noisy image $X_{t}$, and the network predicts the noise $\epsilon_{\theta}$ to restore the original image $\hat{X_{0}}$. Based on the equation
@@
For the first diffusion process, we utilized a weighted combination of mean square error loss (see Eq. 2), identity consistency loss (the same as the Time-step Dependent ID Loss in DCFace [30]), and landmark loss as our first loss function. The identity consistency loss $L_{ID}$ is derived from the cosine similarity between the identity features extracted from the images, while the landmark loss $L_{LM}$ is obtained by calculating the mean squared error of the detected landmark keypoints.
To align with the architecture of CCFExp, we introduced the concept of a secondary loss function to enhance the effectiveness of the model’s training, whose basic composition is consistent with that of the first loss. However, the noisy image $\hat{X_{t}}$ generated from $\hat{X_{0}}$ does not participate in the denoising process after the second diffusion, and $\hat{X_{0}}$ itself is also excluded from the calculation of the second loss. Instead, we calculate the respective loss components using the original image $X_{0}$ before the first diffusion and the restored predicted image $\hat{X_{0}}^{{}^{\prime}}$ after the second denoising. The final loss is the weighted sum of the first loss and the second loss.
Algorithm 1 Training Algorithm of CCFExp
[…]
@@
         Calculate $\mathcal{L}_{lm}=\|D(\hat{x}_{0})-D(x_{0})\|^{2}$
         Calculate initial identity consistency loss
         $\mathcal{L}_{id1}=CS(F(x_{id}),F(\hat{x}_{0}))$
         $\mathcal{L}_{id2}=CS(F(x_{0}),F(\hat{x}_{0}))$
         $\mathcal{L}_{id}=-\frac{t}{T}\mathcal{L}_{id1}-\frac{T-t}{T}\mathcal{L}_{id2}$
         Calculate $\mathcal{L}_{first}=\mathcal{L}_{mse}+\lambda\mathcal{L}_{id}+\gamma\mathcal{L}_{lm}$
@@
         Calculate $\mathcal{L}^{{}^{\prime}}_{mse}=\|\epsilon-\epsilon_{\theta}^{{}^{\prime}}(x_{t}^{{}^{\prime}},x_{id},t)\|^{2}$
@@
         Calculate $\mathcal{L}^{{}^{\prime}}_{lm}=\|D(\hat{x}^{{}^{\prime}}_{0})-D(x_{0})\|^{2}$
         Calculate second identity consistency loss
         $\mathcal{L}^{{}^{\prime}}_{id1}=CS(F(x_{id}),F(\hat{x}^{{}^{\prime}}_{0}))$
         $\mathcal{L}^{{}^{\prime}}_{id2}=CS(F(x_{0}),F(\hat{x}^{{}^{\prime}}_{0}))$
         $\mathcal{L}^{{}^{\prime}}_{id}=-\frac{t}{T}\mathcal{L}^{{}^{\prime}}_{id1}-\frac{T-t}{T}\mathcal{L}^{{}^{\prime}}_{id2}$
         Calculate $\mathcal{L}_{second}=\mathcal{L}^{{}^{\prime}}_{mse}+\lambda\mathcal{L}^{{}^{\prime}}_{id}+\gamma\mathcal{L}^{{}^{\prime}}_{lm}$

### v3

# CFCPalsy: Facial Image Synthesis with Cross-Fusion Cycle Diffusion Model for Facial Paralysis Individuals

@@ Abstract
Currently, the diagnosis of facial paralysis remains a challenging task, often relying heavily on the subjective judgment and experience of clinicians, which can introduce variability and uncertainty in the assessment process. One promising application in real-life situations is the automatic estimation of facial paralysis. However, the scarcity of facial paralysis datasets limits the development of robust machine learning models for automated diagnosis and therapeutic interventions. To this end, this study aims to synthesize a high-quality facial paralysis dataset to address this gap, enabling more accurate and efficient algorithm training. Specifically, a novel Cross-Fusion Cycle Palsy Expression Generative Model (CFCPalsy) based on the diffusion model is proposed to combine different features of facial information and enhance the visual details of facial appearance and texture in facial regions, thus creating synthetic facial images that accurately represent various degrees and types of facial paralysis. We have qualitatively and quantitatively evaluated the proposed method on the commonly used public clinical datasets of facial paralysis to demonstrate its effectiveness. Experimental results indicate that the proposed method surpasses state-of-the-art methods, generating more realistic facial images and maintaining identity consistency. Code Link
@@ 1 Introduction
Facial palsy, which can result in temporary or permanent weakness or paralysis of one side of the face, affects approximately 23 out of every 100,000 people annually [16]. Beyond the physical discomfort, patients with facial palsy often experience significant psychological distress [22]. Accurate and timely diagnosis is crucial for effective treatment; however, the variability in clinical presentation poses significant challenges, and the severity estimation of facial palsy remains a largely subjective process [16].
@@
In recent years, machine learning (ML) has shown promise in aiding the diagnosis of facial paralysis by analyzing facial features and expressions [63, 1, 15, 16]. However, the development of reliable ML models is hindered by the scarcity of comprehensive, annotated datasets that accurately represent the diversity and severity of facial paralysis. Obtaining high-quality facial images of facial palsy patients is a significant challenge due to several factors, including the need to protect patient privacy and the inherent sensitivity of medical data. Additionally, the complexity of the condition and the high labor costs involved make it challenging to collect a diverse and representative dataset. Existing datasets [65, 18, 25] are often limited in size, scope, and variability, restricting the generalization ability and performance of ML algorithms.
@@
Several studies have explored the use of generative adversarial networks (GANs) [17] to generate facial paralysis face images [67, 51, 50]. Although [67] successfully utilizes GANs to simulate facial expressions by reconstructing facial images from extracted boundary information, the generated images often lack the fine details necessary for accurate medical diagnosis, especially when there is a significant disparity between the facial features of the patient and the face to be combined with the palsy features. Building on this, another study [51] introduces facial shape normalization to address the unnatural appearance of images resulting from shape differences between the patient and the model. While this method improves image realism, it still struggles with capturing complex facial expressions and fine details, limiting its effectiveness in more nuanced diagnostic scenarios. In addition, to leverages data augmentation, GANs are also used to improve the performance of convolutional neural networks (CNNs) in grading facial palsy [50]. This approach synthesizes additional face images with varying degrees of facial palsy to combat overfitting and enhance the model’s generalization capability. However, the synthesized images in this study primarily focus on only five categories without adequately capturing the full range of facial expressions and subtle asymmetries inherent in real-world cases.
@@
In this paper, we aim to address the critical need for a comprehensive facial paralysis dataset by synthesizing a high-quality dataset. The generator is designed to capture a wide range of facial paralysis conditions and able to generate face images with varying degrees of facial paralysis based on a given facial identity image, providing a valuable resource for training ML models with the potential to improve diagnostic accuracy and clinical outcomes.
@@
To this end, we propose a Cross-Fusion Cycle Palsy Expression Generative Model (CFCPalsy). CFCPalsy is a unique triple condition image generation model, conditioned on identity information of a given ID image, expression information and landmark information of a given style image. It could generate a realistic image of the face in the ID image exhibiting the facial palsy characteristics of the face in the style image. CFCPalsy employs landmark features and the corresponding loss to ensure that the model could fully capture the nuanced facial deformations, particularly the detailed changes in facial features such as eyes, mouth, and overall asymmetry. When the model is fed multiple features simultaneously, it often encounters challenges such as feature redundancy and conflicting information, which can lead to degraded performance [36]. The redundancy among features can cause the model to overweight certain aspects, while conflicting signals may lead to inconsistent learning and prediction. To address these issues, CFCPalsy utilizes a novel cross-fusion feature integration method. This approach effectively harmonizes the information from multiple features, enabling the model to better leverage the complementary aspects of each feature, thereby significantly improving the overall performance and robustness of the model. Training diffusion models [20, 54] on limited datasets often leads to inefficiencies in learning and suboptimal model performance due to the scarcity of data. To address these challenges, we focused on improving the training efficiency and overall performance of diffusion models under constrained data conditions by modifying the training strategy. Our approach optimizes the learning process and proposes the cycle diffusion framework enabling the model to achieve superior results with a limited dataset.
@@
  • Our approach introduces a novel application of generative models for facial paralysis synthesis, achieving state-of-the-art results and addressing the scarcity of medical datasets, thereby facilitating future research in automated facial analysis.
Figure 2: An overview of the forward process of CFCPalsy. a) The architecture of our feature extractors. We train the model using different images of the same patient, with one serving as the identity image $X_{ID}$ and the other as the facial palsy expression image $X_{0}$. The identity model extracts identity features from the identity image. The expression module (including $E_{Face}$ and $E_{ID}$) and $E_{LM}$ extract expression and landmark features from the facial palsy expression image, respectively. b) An illustration of the principle of the cross feature fusion strategy. We employ a method similar to cross-attention to facilitate information exchange among the three conditional features. After this interaction, the features are concatenated for further processing. c) A diagram of the noise predicting network. CFCPalsy utilizes a classic U-Net architecture combined with residual connections and cross-attention mechanisms to accurately predict the noise.
@@ 3 Method
In this section, we present our Cross-Fusion Cycle Palsy Expression Model (CFCPalsy) designed to synthesize high-quality facial paralysis images (see Fig. 2). First, we mention the necessary preliminary concepts about the diffusion models (sec 3.1). Next, we introduce the various feature extractors used in CFCPalsy which enhance the diffusion model’s ability to capture more facial details by accurately extracting multiple features (sec 3.2). In sec 3.3, we propose the cross feature fusion, which mitigates the model’s over-reliance on individual features and enhances its overall learning capability across all features by exchanging query vectors formed from different features. Finally, we devise the cycle training strategy that improves the training efficiency and performance of the diffusion model on limited datasets by modifying the training process and introducing a secondary loss function (sec 3.4).
@@ 3.1 Preliminary
$L(\theta)=\mathbb{E}_{t,X_{t},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(X_{t},t)\|^{2}\right],$ (2)
@@ 3.2 Feature Extraction Module
ID Extraction Module
To accurately capture identity information from $X_{ID}$, we use the IR-50 [10], a typical CNN, as the identity feature extractor. We utilize both the output features $f_{ID}\in\mathbb{R}^{512}$ from the final layer and the intermediate layer features $I_{ID}\in\mathbb{R}^{512\times 7\times 7}$of the model to enhance the quality of image generation. For $I_{ID}$, we first flatten and rearrange their dimensions to obtain $F_{ID}\in\mathbb{R}^{49\times 512}$ that facilitates subsequent processing. Then, we concatenate the two sets of features and add positional embeddings $P_{emb}$. In summary, the output of the identity feature extraction module $M_{ID}$ is formulated as follows:
$M_{ID}(X_{ID})=Concat[f_{ID},Flatten(I_{ID})]+P_{emb}.$ (3)
@@
Expression Extraction Module
CFCPalsy employs the deviation framework of DLN [71] to accurately capture the expression features. This module $M_{Exp}$ consists of two models: a facial feature extraction model $E_{Face}$ and an identity feature extraction model $E_{ID}$, sharing the same network structure. Based on the assumption that facial features can be divided into identity features and expression features, we use two separate models to extract expression and identity features from the same facial paralysis image $X_{0}$. By subtracting the identity features from the facial features, we obtain expression features that are invariant to identity. The subtracted features are then passed through a convolution operation for channel alignment. Specifically,
$M_{Exp}(X_{0})=Conv[E_{Face}(X_{0})-E_{ID}(X_{0})].$ (4)
@@
Facial landmarks refer to key points on a face such as the corners of the eyes, the tip of the nose, and the edges of the lips that are used to represent and understand facial geometry, which could enhance facial generation by imposing constraints on critical facial regions. For the first time, we integrate facial landmark features with diffusion models, coupled with a specifically designed loss function, to generate realistic and detailed facial images of patients with facial palsy. In our CFCPalsy, we employ the MobileFaceNet [7] as the landmark detector. We remove its output layer to use it as a landmark feature extractor $E_{LM}$, which outputs a feature $F_{LM}\in\mathbb{R}^{49\times 512}$.
@@ 3.3 Cross-Fusion Module
The attention mechanism [62] has become a fundamental component in many state-of-the-art models, particularly for tasks requiring the integration of information from different sources or across different parts of a data sequence. At its core, the attention mechanism allows a model to weigh the importance of different elements within a sequence, dynamically focusing on the most relevant parts to improve decision-making or feature extraction. The cross-attention mechanism extends this concept to operate between two distinct sequences. This facilitates the fusion of information from different sources, making it particularly useful for tasks that require integrating features from multiple modalities or input streams.
We aim for the model to fully utilize the information from identity, expression, and landmarks when predicting noise. Although directly concatenating the features from these three sources and injecting them into the model is a feasible approach, it does not yield satisfactory results. The generated images either lose the identity information from identity images $X_{ID}$ or fail to retain the facial expression details from original images $X_{0}$. Building on the concept of the cross-attention mechanism, we perform mutual interaction among multiple features, where the query sequence from one feature guides the extraction of information from the key-value sequence of another feature.
The features from the extractors are first mapped into the query, key, and value vector spaces separately through a QKV encoder. Specifically,
@@
where $F$ indicates the three types of features extracted from the former modules, and $W^{Q},W^{K},W^{V}$ represent the model’s learnable parameter matrixes. Similar to the multi-head attention mechanism [62], we utilize multiple sets of transformation matrices to obtain multiple vector spaces, aiming to capture different interaction patterns among features.
Subsequently, the identity features, expression features, and landmark features each have their own corresponding QKV vector spaces. Next, we perform vector interactions. Specifically,
$\text{Attention}_{ID}=\text{Softmax}\left(\frac{Q_{LM}K_{ID}^{T}}{\sqrt{d_{k}}}\right)V_{ID},$ (6)
$\text{Attention}_{Exp}=\text{Softmax}\left(\frac{Q_{ID}K_{Exp}^{T}}{\sqrt{d_{k}}}\right)V_{Exp},$ (7)
$\text{Attention}_{LM}=\text{Softmax}\left(\frac{Q_{Exp}K_{LM}^{T}}{\sqrt{d_{k}}}\right)V_{LM},$ (8)
where $d_{k}$ is a scaling factor.
Figure 3: An illustration of the cycle training strategy of CFCPalsy. Each training data undergoes two diffusion processes, where the plus sign indicates the noise addition process.
@@ 3.4 Cycle Diffusion Strategy
Building upon the training framework of the original diffusion model, we incorporated an additional step of secondary diffusion. Algorithm 1 shows the training procedure of CFCPalsy. After the original diffusion process, where the input image $X_{0}$ is noised to generate a noisy image $X_{t}$, and the network predicts the noise $\epsilon_{\theta}$ to restore the original image $\hat{X_{0}}$. Based on the equation
@@
For the first diffusion process, we utilized a weighted combination of mean square error loss (see Eq. 2), identity consistency loss $L_{ID}$ (the same as the Time-step Dependent ID Loss in DCFace [31]), and landmark loss $L_{LM}$ as our first loss function. The identity consistency loss $L_{ID}$ is derived from the cosine similarity between the identity features extracted from the images, while the landmark loss $L_{LM}$ is obtained by calculating the mean squared error of the detected landmark keypoints:
$L_{first}=L(\theta)+\lambda L_{ID}(X_{ID},\hat{X_{0}})+\gamma L_{LM}(X_{0},\hat{X_{0}}),$ (10)
To align with the architecture of CFCPalsy, we introduced the concept of a secondary loss function to enhance the effectiveness of the model’s training, whose basic composition is consistent with that of the first loss. However, the noisy image $\hat{X_{t}}$ generated from $\hat{X_{0}}$ does not participate in the denoising process after the second diffusion, and $\hat{X_{0}}$ itself is also excluded from the calculation of the second loss. Instead, we calculate the respective loss components using the original image $X_{0}$ before the first diffusion and the restored predicted image $\hat{X_{0}}^{{}^{\prime}}$ after the second denoising. Specifically,
$L_{second}=L(\theta^{\prime})+\lambda L_{ID}(X_{ID},\hat{X_{0}^{\prime}})+\gamma L_{LM}(X_{0},\hat{X_{0}^{\prime}}),$ (11)
The final loss is the weighted sum of the first loss and the second loss.
Algorithm 1 Training Algorithm of CFCPalsy
@@
         Calculate $\mathcal{L}_{first}$
@@
         Calculate $\mathcal{L}_{second}$